OpenStudy (lxelle):

solve the inequality 2-3x < |x-3|.

OpenStudy (lxelle):

So 8x^2 -6x-5<0 x=-1/2 and x=5/4 But why is th3 answer only x=-1/2??

OpenStudy (anonymous):

Well, you shouldn't have equals given that it's an inequality

OpenStudy (anonymous):

However, you should have: \[ 2-3x<x-3 \]As well as \[ 2-3x<-(x-3) \]And you simplify them.

OpenStudy (lxelle):

But theres modulus. I can't put the negative there.

OpenStudy (anonymous):

You solve modulus by breaking it down into both possible cases: the positive case and the negative case.

OpenStudy (lxelle):

Okay I did (2-3x)^2 < (x-3)^2

OpenStudy (lxelle):

Then?

OpenStudy (anonymous):

Using that method is a bit problematic

OpenStudy (lxelle):

It requires to use this method.

OpenStudy (anonymous):

Squaring both sides of the inequality doesn't always work that way

OpenStudy (lxelle):

The answerscheme says that gives a mark.

OpenStudy (lxelle):

And afterall I need to find wheather its less or more than

OpenStudy (anonymous):

Positive case:\[ 2-3x<x-3\implies 5<4x \implies \frac 54 <x \]Negative case:\[ 2-3x<-(x-3)\implies -2+3x>x-3\implies 2x>-1 \implies x>-\frac 12 \]

OpenStudy (anonymous):

Now \(-1/2 < 5/4\).

OpenStudy (lxelle):

What if i wanna use my method? Is there any way to find ou t whether it's less or ,ore?

OpenStudy (anonymous):

@wio 's method is the easiest and simplest method.

OpenStudy (lxelle):

I,m just asking.

OpenStudy (anonymous):

For either to be true: \(-1/2<x\). For both to be true: \(5/4<x\). I believe that with modulus, we use the either restriction.

OpenStudy (lxelle):

I actually tried proving x<5/4 in the equation and it can't seem To be proven

OpenStudy (anonymous):

Here is the problem with your method.

OpenStudy (anonymous):

\[ -10<|5| \]Square both sides: \[ 100 < 25 \]This is false.

OpenStudy (lxelle):

Is your answer x>-1/2???

OpenStudy (lxelle):

????

OpenStudy (lxelle):

@Kainui

OpenStudy (adi3):

i think it should be 2-3x < x+3

OpenStudy (adi3):

bc its a absolute value the sign changes

OpenStudy (adi3):

@LXelle

OpenStudy (lxelle):

Uh no. Modulus means sqyare both sides

OpenStudy (adi3):

ohh yeah i forgot